OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

OpenStudy (johnweldon1993):

we go from standard form to slope intercept form...y=mx + b 6x - 3y = 12 -3y = 12 - 6x y = 12/-3 + 6x/3 y = -4 + 2x I get slope is 2 and y intercept is (0,-4)